A rectangle has an area of 60 square meters and a perimeter of 38 meters. find the dimensions of the rectangle.
Answer by checkley77(12844)

XY=60 OR X=60/Y
2X+2Y=38
2(60/Y)+2Y=38
120/Y+2Y=38
(120+2Y*Y)/Y=38
(120+2Y^2)/Y=38
120+2Y^2=38Y
2Y^2-38Y+120=0
2(Y^2-19Y+60)=0
2(Y-15)(Y-4)=0
Y-15=0
Y=15 ANS
Y-4=0
Y=4 ANS.
X*4=60
X=60/4
X=15 ANS
X*15=60
X=60/15
X=4 ANS.
PROOF:
2*15+2*4=38
30+8=38
38=38